OpenStudy (anonymous):

p=sina+sin2a+sin3a+........+sinna

OpenStudy (ybarrap):

This is a geometric series, with \( a_n = sin(a) \): http://en.wikipedia.org/wiki/Geometric_series The solution you get will be valid for all \(a\) not a multiple of \(2\pi\), including 0. Let me know if you have any questions.

OpenStudy (tkhunny):

Is that second term \(\sin(2a)\;or\;\sin^{2}(a)\)?